# Polylogarithmic motivic Chabauty–Kim for $\mathbb{P}^{1}\setminus\{0,1,\infty\}$: the geometric step via resultants

### v2

@@ Abstract.
Given a finite set $S$ of distinct primes, we propose a method to construct polylogarithmic motivic Chabauty–Kim functions for $\mathbb{P}^{1}\setminus\{0,1,\infty\}$ using resultants. For a prime $p\not\in S$, the vanishing loci of the images of such functions under the $p$-adic period map contain the solutions of the $S$-unit equation. In the case $|S|=2$, we explicitly construct a non-trivial motivic Chabauty–Kim function in depth 6 of degree 18, and prove that there do not exist any other Chabauty–Kim functions with smaller depth and degree. The method, inspired by work of Dan-Cohen and the first author, enhances the geometric step algorithm developed by Corwin and Dan-Cohen, providing a more efficient approach.
@@ 1. Introduction
  • Geometric step: find non-trivial polylogarithmic motivic Chabauty–Kim functions in depth $\leq d$. These functions (Definition 3.3) are elements of a weighted polynomial ring $\mathcal{O}(U_{S})[\log^{\mathfrak{u}},\operatorname{Li}^{\mathfrak{u}}_{1},\ldots,\operatorname{Li}^{\mathfrak{u}}_{d}]$, which vanish under a certain cocyle map $\theta^{\#}$ (Definition 4.1). Here:
    • the variables are the unipotent de Rham polylogarithms $\log^{\mathfrak{u}}$ and $\operatorname{Li}_{n}^{\mathfrak{u}}$ for $n\geq 1$ (see Section 2.3);
    • the coefficient ring $\mathcal{O}(U_{S})$ is the coordinate ring of the pro-unipotent mixed Tate fundamental group over $\mathbb{Z}_{S}$ (see Section 2.2);
@@
  • Arithmetic step: determine the images of the functions found in the geometric step under the $p$-adic period map, in order to obtain $p$-adic Chabauty–Kim functions. [^1] Guaranteed to be non-trivial if one assumes the $p$-adic Period Conjecture [CDC1, Conj. 2.25]. This step depends on the prime $p$ and the specific primes contained in $S$.
@@ 3. $(iii)$
    There exists a non-trivial motivic Chabauty–Kim function $F^{|2|}_{6,18}$ of depth $6$ and degree $18$ (see Definition 6.9). Moreover, any other motivic Chabauty–Kim function of depth less than $19$ and degree $18$ is a multiple of $F^{|2|}_{6,18}$ by an element of $\mathcal{O}(U_{S})$.
This result constitutes the first progress towards explicit Chabauty–Kim in depth $>4$ for $|S|\geq 2$.
@@
The cocycle map $\theta^{\#}$ can be viewed as a morphism of graded vector spaces, and as the depth $d$ and the degree $v$ grow, the dimension of the relevant graded piece of the domain grows faster than that of the corresponding graded piece of the codomain (Proposition 5.2); we tabulate these dimensions in Table 1. In particular, when $d$ and $v$ are large enough, the existence of a non-trivial motivic Chabauty–Kim function is guaranteed by a simple dimension argument. However, the non-trivial motivic Chabauty–Kim function $F^{|2|}_{6,18}$ of Theorem 1.1 does not fit into this regime: the depth $6$, degree $18$ part of the domain has dimension $996$, while the corresponding part of the codomain has dimension $4183$. Hence, the mere existence of $F^{|2|}_{6,18}$ is surprising! At least when $|S|=2$, our construction can produce multiple distinct non-trivial motivic Chabauty–Kim functions whose existence cannot be explained by naive dimension estimates. It would be interesting to see what circumstances could give an a priori justification for their existence. These observations offer hints of new phenomena in Chabauty–Kim theory that are not visible when $|S|=1$ (see Section 5.2.1).
@@ 1.1. Polylogarithmic motivic Chabauty–Kim
We refer to Section 3.3 for a detailed description of this diagram. In this context:
  • $\Pi^{\operatorname{PL}}$ denotes the polylogarithmic pro-unipotent de Rham fundamental group of $X$ (see Section 2.3);
  • the Selmer scheme is defined as $\operatorname{Sel}_{S}^{\operatorname{PL}}=\operatorname{Hom}_{\mathbb{G}_{m}}(U_{S},\Pi^{\operatorname{PL}})$, where $U_{S}$ denotes the pro-unipotent mixed Tate fundamental group over $\mathbb{Z}_{S}$ (see Section 2.2);
@@
[…]
@@ 1.2. Methods
When $|S|=1$, Corwin and Dan-Cohen compute the kernel of $\theta^{\#}_{d,v}$ for small $d$ and $v$ by hand, with minimal difficulty. However, this case is deceptively simple: we show in Remark 5.8 that for $|S|=2$, even when one knows that $\ker(\theta^{\#}_{6,18})$ should be non-trivial, computing its kernel directly amounts to computing the kernel of a $4183\times 996$ matrix over a polynomial algebra in $30$ variables, which is infeasible even for a supercomputer.
To prove Theorem 1.1, we explicitly compute this kernel using a two-faceted approach. First, we develop an algorithm to prove that $\ker(\theta^{\#}_{d,v})=0$ whenever $d,v\leq 17$. Second, via the method of resultants, we explicitly construct an element of $\ker(\theta^{\#}_{6,18})$.
@@ 1.2.1. An algorithm for bounding the number of motivic Chabauty–Kim functions
We use this observation to formulate Algorithm 5.6, which takes as input a tuple of positive integers $(s,d,v)$, where $s=|S|$, and outputs the dimension of the kernel of $\theta^{\#}_{d,v}$, evaluated at a random integer vector $x\in\mathbb{Z}^{n}$. The output is therefore an upper bound (and likely also a lower bound – see Remark 5.5) for the dimension of the space of polylogarithmic motivic Chabauty–Kim functions in depth $d$ and degree $v$ for $s$ primes (Proposition 5.4).
We have implemented this algorithm in SageMath [sagemath]; the code can be viewed at our GitHub repository. The output of the algorithm for $(2,17,17)$ is 00, proving Theorem 1.1$(i)$ (Theorem 5.7).
However, the output of the algorithm for $(2,17,18)$, and even $(2,6,18)$, is $1$, which indicates that it is extremely likely that a non-zero polylogarithmic motivic Chabauty–Kim function exists in depth $6$ and degree $18$. Granted the existence of such a function, the output proves Theorem 1.1$(iii)$. In order to prove Theorem 1.1, all that remains is therefore to prove the existence of such a function.
@@ 1.2.2. Constructing motivic Chabauty–Kim functions via resultants
For any $|S|$, we propose a resultant method that effectively constructs polylogarithmic motivic Chabauty–Kim functions (Section 7). The method is inspired by the work [DCJ] on the surface $M_{0,5}$ (the moduli space of genus 00 curves with $5$ marked points). At this level of generality, we cannot guarantee the non-triviality of the resulting functions. However, when $|S|=2$, we construct a function of depth $6$ and degree $18$ and prove its non-triviality (Theorem 6.10). We now briefly sketch the construction in the case $|S|=2$ (see Section 6 for the detailed construction).
Constructing a polylogarithmic motivic Chabauty–Kim function requires finding an element in the kernel of the map (1.3). As explained in Section 1.2.1, after restricting to a fixed depth $d$ and degree $v$, the map (1.4) can be expressed as a matrix with coefficients in the shuffle algebra $\mathcal{O}(U_{S})_{\leq d}$. It is however computationally infeasible to compute its null space, even though we can implement the matrix in SageMath. In the case where $|S|=2$ and $d$ is even, the images of $\log^{\mathfrak{u}},\operatorname{Li}_{1}^{\mathfrak{u}},\ldots,\operatorname{Li}_{d}^{\mathfrak{u}}$ under (1.3) take values in a polynomial algebra with $24+(d-1)/$ 2 variables $x_{1},y_{1},x_{2},y_{2},z_{3},z_{5},\ldots,z_{d-1}$ and coefficients in $\mathcal{O}(U_{S})_{\leq d}$. Due to the particular shape of the map (1.3), it is possible to express $z_{k}$ in terms of images under (1.3) and the variables $x_{1},y_{1},x_{2},y_{2},$ and $z_{k^{\prime}}$ for $k^{\prime}<k$. From the shape of $\theta^{\#}(\log^{\mathfrak{u}}),\theta^{\#}(\operatorname{Li}_{1}^{\mathfrak{u}})$, and $\theta^{\#}(\operatorname{Li}_{2}^{\mathfrak{u}})$, the variables $x_{1},y_{1},x_{2},y_{2}$ can in turn be expressed in terms of images under (1.2) and the single variable $x_{1}$. From these observations, it is possible to derive a polynomial expression $P_{d}(X)\in\mathrm{Im}(\theta^{\#})[X]$ with the property that $P_{d}(x_{1})=0$. The construction works whenever $d\geq 4$. Write $P_{d}(X)=\theta^{\#}(\nu_{d}(X))$ for some $\nu_{d}(X)\in(\mathcal{O}(U_{S})_{\leq d}[\log^{\mathfrak{u}},\operatorname{Li}_{1}^{\mathfrak{u}},\ldots,\operatorname{Li}_{d}^{\mathfrak{u}}])[X]$. Then
@@
is a non-trivial Chabauty–Kim function of depth $6$ and degree $18$ (Theorem 6.10). It is worth pointing out that the factor $\operatorname{Li}_{2}^{\mathfrak{u}}-\frac{1}{2}\log^{\mathfrak{u}}\operatorname{Li}_{1}^{\mathfrak{u}}$ is precisely the generator of {},$\mathcal{I}^{\operatorname{PL}}_{\{\ell\},2}$ studied in [DCW2], where $\ell$ denotes any prime. We note however that it is not an element of $\mathcal{I}^{\operatorname{PL}}_{S,2}$ for $|S|=2$.
@@ 1.3. Prior works
The present work is the first to tackle explicit Chabauty–Kim in depth $>4$ for $|S|\geq 2$. In this section, we give an overview of prior works on the subject.
The map loc was first made explicit in [DCW1] in depth $2$. In [BDCKW], the case $S=\emptyset$ in depths $1$ and $2$ was studied. In [DCW1], the equality $X(\mathbb{Z}_{S})=(X(\mathbb{Z}_{p})^{\mathrm{PL}}_{S,d})^{S_{3}}$ was verified for $S=\{2\}$ and $d=2$ with $p=3,5,7$, but turned out to fail with $p=11$. In [DCW2], Dan-Cohen and Wewers studied the case $S=\{2\}$ in depth $4$ and verified the equality $X(\mathbb{Z}_{S})=(X(\mathbb{Z}_{p})^{\mathrm{PL}}_{S,d})^{S_{3}}$ with $p=3,5,\ldots,29$ using two explicit $p$-adic Chabauty–Kim functions in depths $2$ and $4$ respectively [DCW2, Thm. 1.17.1]. The motivic framework of the method was further developed in [DC] resulting in an algorithm for determining $X(\mathbb{Z}_{S})$ whose halting is conditional on a list of conjectures. The polylogarithmic motivic Chabauty–Kim method for $X$ was fully developed in [CDC1], including the separation into the geometric and the arithmetic step. The geometric step was solved for $|S|=1$ in depth $4$, and the equality $X(\mathbb{Z}_{S})=(X(\mathbb{Z}_{p})^{\mathrm{PL}}_{S,d})^{S_{3}}$ was verified for $S=\{3\}$ in depth $4$ with $p=5,7$. It is also in [CDC1] that the necessity of symmetrising the polylogarithmic Chabauty–Kim locus via the $S_{3}$-action was discovered. In [CDC2], an algorithm that determines the set $X(\mathbb{Z}_{S})$ based on the method of [CDC1] is described, whose halting depends on a list of conjectures. Part of the work of the present paper can be seen as making their algorithm for the geometric step effective, a step that was deceptively perceived in [CDC1, CDC2] as simple. In related work, motivic Chabauty–Kim was recently developed for the surface $M_{0,5}$ in [DCJ] and was worked out explicitly for $S=\{2,3\}$ in depth $4$. A certain weight advantage encountered for $M_{0,5}$ allowed Dan-Cohen and the first author to work in depth $4$, a depth that is too small to handle $X$ when $|S|=2$. Their calculations solve in particular the arithmetic step for $X$ in depth $4$ for $S=\{2,3\}$. Recently, the case $S=\{2,3\}$ for $X$ was tackled in [martin] using a different method, namely the refined Chabauty–Kim method first developed in [bettsdogra]. This refined version of the method allows Lüdtke to work in depth $4$, resulting in new evidence for the refined Kim’s conjecture in this setting. The previous work [BB+] concerned the refined Chabauty–Kim method for $|S|=2$ in depth $2$. The recent work [BKL24] proves Kim’s conjecture for $S=\emptyset$ and the refined Kim’s conjecture for $S=\{2\}$, for all odd primes $p$.
@@ 1.4. Outline
Section 2 contains preliminaries on mixed Tate motives and motivic fundamental groups. Section 3 gives an overview of the polylogarithmic motivic Chabauty–Kim method. Section 4 gives a description of the evaluation map (1.2) in terms of non-canonical coordinates of the Selmer scheme. This marks the end of the background material. Section 5 contains the upper bound algorithm and the proofs of Theorem 1.1$(i)$, $(iii)$. Finally, in Section 6, we describe the resultant method, which culminates in the proof of Theorem 1.1. Section 7 contains a brief discussion of possible generalisations of the method to the case $|S|>2$.

### v3

@@ Abstract.
Given a finite set $S$ of primes, we propose a method to construct polylogarithmic motivic Chabauty–Kim functions for $\mathbb{P}^{1}\setminus\{0,1,\infty\}$ using resultants. For a prime $p\not\in S$, the vanishing loci of the images of such functions under the $p$-adic period map contain the solutions of the $S$-unit equation. In the case $|S|=2$, we explicitly construct a non-trivial motivic Chabauty–Kim function in depth 6 of degree 18, and prove that there do not exist any other Chabauty–Kim functions with smaller depth and degree. The method, inspired by work of Dan-Cohen and the first author, enhances the geometric step algorithm developed by Corwin and Dan-Cohen, providing a more efficient approach.
@@ 1. Introduction
  • Geometric step: find non-trivial polylogarithmic motivic Chabauty–Kim functions in depth $\leq d$. These functions (Definition˜3.3) are elements of a weighted polynomial ring $\mathcal{O}(U_{S})[\log^{\mathfrak{u}},\operatorname{Li}^{\mathfrak{u}}_{1},\ldots,\operatorname{Li}^{\mathfrak{u}}_{d}]$, which vanish under a certain cocyle map $\theta^{\#}$ (Definition˜4.1). Here:
    • the variables are the unipotent de Rham polylogarithms $\log^{\mathfrak{u}}$ and $\operatorname{Li}_{n}^{\mathfrak{u}}$ for $n\geq 1$ (see Section˜2.3);
    • the coefficient ring $\mathcal{O}(U_{S})$ is the coordinate ring of the pro-unipotent mixed Tate fundamental group over $\mathbb{Z}_{S}$ (see Section˜2.2);
@@
  • Arithmetic step: determine the images of the functions found in the geometric step under the $p$-adic period map, in order to obtain $p$-adic Chabauty–Kim functions.  [^1] Can be checked to be non-trivial in practice and are guaranteed to be non-trivial if one assumes the $p$-adic Period Conjecture [CDC1, Conj. 2.25]. This step depends on the prime $p$ and the specific primes contained in $S$.
@@ 3. $(iii)$
    There exists a non-trivial motivic Chabauty–Kim function $F^{|2|}_{6,18}$ of depth $6$ and degree $18$ (see Definition˜6.9). Moreover, any other motivic Chabauty–Kim function of depth less than $19$ and degree $18$ is a multiple of $F^{|2|}_{6,18}$ by an element of $\mathcal{O}(U_{S})$.
This result constitutes the first progress towards explicit Chabauty–Kim in depth >4>4 for $|S|\geq 2$.
@@
The cocycle map $\theta^{\#}$ can be viewed as a morphism of graded vector spaces, and as the depth $d$ and the degree $v$ grow, the dimension of the relevant graded piece of the domain grows faster than that of the corresponding graded piece of the codomain (Proposition˜5.3); we tabulate these dimensions in Table˜1. In particular, when $d$ and $v$ are large enough, the existence of a non-trivial motivic Chabauty–Kim function is guaranteed by a simple dimension argument. However, the non-trivial motivic Chabauty–Kim function $F^{|2|}_{6,18}$ of Theorem˜1.1 does not fit into this regime: the depth $6$, degree $18$ part of the domain has dimension $996$, while the corresponding part of the codomain has dimension $4183$. Hence, the mere existence of $F^{|2|}_{6,18}$ is surprising! At least when $|S|=2$, our construction can produce multiple distinct non-trivial motivic Chabauty–Kim functions whose existence cannot be explained by naive dimension estimates. It would be interesting to see what circumstances could give an a priori justification for their existence. These observations offer hints of new phenomena in Chabauty–Kim theory that are not visible when $|S|=1$ (see Section 5.2.1).
@@ 1.1. Polylogarithmic motivic Chabauty–Kim
We refer to Section˜3.3 for a detailed description of this diagram. In this context:
  • $\Pi^{\operatorname{PL}}$ denotes the polylogarithmic pro-unipotent de Rham fundamental group of $X$ (see Section˜2.3);
  • the Selmer scheme is defined as $\operatorname{Sel}_{S}^{\operatorname{PL}}=\operatorname{Hom}_{\mathbb{G}_{m}}(U_{S},\Pi^{\operatorname{PL}})$, where $U_{S}$ denotes the pro-unipotent mixed Tate fundamental group over $\mathbb{Z}_{S}$ (see Section˜2.2);
@@
$f|_{X(\mathbb{Z}_{p})}\colon X(\mathbb{Z}_{p})\xrightarrow{\kappa}\Pi^{\operatorname{PL}}(\mathbb{Q}_{p})\xrightarrow{f_{\mathbb{Q}_{p}}}\mathbb{Q}_{p}.$
@@ 1.2. Methods
[…]
@@ 1.2.1. An algorithm for bounding the number of motivic Chabauty–Kim functions
We use this observation to formulate Algorithm˜5.7, which takes as input a tuple of positive integers $(s,d,v)$, where $s=|S|$, and outputs the dimension of the kernel of $\theta^{\#}_{d,v}$, evaluated at a random integer vector $x\in\mathbb{Z}^{n}$. The output is therefore an upper bound (and likely also a lower bound – see Remark 5.6) for the dimension of the space of polylogarithmic motivic Chabauty–Kim functions in depth $d$ and degree $v$ for $s$ primes (Proposition˜5.5).
We have implemented this algorithm in SageMath [sagemath]; the code can be viewed at our GitHub repository. The output of the algorithm for $(2,17,17)$ is 0, proving Theorem˜1.1(i)(i) (Theorem˜5.8).
However, the output of the algorithm for $(2,17,18)$, and even $(2,6,18)$, is $1$, which indicates that it is extremely likely that a non-zero polylogarithmic motivic Chabauty–Kim function exists in depth $6$ and degree $18$. Granted the existence of such a function, the output proves Theorem˜1.1(iii)(iii). In order to prove Theorem˜1.1, all that remains is therefore to prove the existence of such a function.
@@ 1.2.2. Constructing motivic Chabauty–Kim functions via resultants
For any $|S|$, we propose a resultant method that effectively constructs polylogarithmic motivic Chabauty–Kim functions (Section˜7). The method is inspired by the work [DCJ] on the surface $M_{0,5}$ (the moduli space of genus 0 curves with $5$ marked points). At this level of generality, we cannot guarantee the non-triviality of the resulting functions. However, when $|S|=2$, we construct a function of depth $6$ and degree $18$ and prove its non-triviality (Theorem˜6.10). We now briefly sketch the construction in the case $|S|=2$ (see Section˜6 for the detailed construction).
Constructing a polylogarithmic motivic Chabauty–Kim function requires finding an element in the kernel of the map (1.3). As explained in Section˜1.2.1, after restricting to a fixed depth $d$ and degree $v$, the map (1.4) can be expressed as a matrix with coefficients in the shuffle algebra $\mathcal{O}(U_{S})_{\leq d}$. It is however computationally infeasible to compute its null space, even though we can implement the matrix in SageMath. In the case where $|S|=2$ and $d$ is even, the images of $\log^{\mathfrak{u}},\operatorname{Li}_{1}^{\mathfrak{u}},\ldots,\operatorname{Li}_{d}^{\mathfrak{u}}$ under (1.3) take values in a polynomial algebra with 4+(d1)/24+(d-1)/2 variables $x_{1},y_{1},x_{2},y_{2},z_{3},z_{5},\ldots,z_{d-1}$ and coefficients in $\mathcal{O}(U_{S})_{\leq d}$. Due to the particular shape of the map (1.3), it is possible to express $z_{k}$ in terms of images under (1.3) and the variables $x_{1},y_{1},x_{2},y_{2},$ and $z_{k^{\prime}}$ for $k^{\prime}<k$. From the shape of $\theta^{\#}(\log^{\mathfrak{u}}),\theta^{\#}(\operatorname{Li}_{1}^{\mathfrak{u}})$, and $\theta^{\#}(\operatorname{Li}_{2}^{\mathfrak{u}})$, the variables $x_{1},y_{1},x_{2},y_{2}$ can in turn be expressed in terms of images under (1.2) and the single variable $x_{1}$. From these observations, it is possible to derive a polynomial expression $P_{d}(X)\in\mathrm{Im}(\theta^{\#})[X]$ with the property that $P_{d}(x_{1})=0$. The construction works whenever $d\geq 4$. Write $P_{d}(X)=\theta^{\#}(\nu_{d}(X))$ for some $\nu_{d}(X)\in(\mathcal{O}(U_{S})_{\leq d}[\log^{\mathfrak{u}},\operatorname{Li}_{1}^{\mathfrak{u}},\ldots,\operatorname{Li}_{d}^{\mathfrak{u}}])[X]$. Then
@@
is a non-trivial Chabauty–Kim function of depth $6$ and degree $18$ (Theorem˜6.10). It is worth pointing out that the factor $\operatorname{Li}_{2}^{\mathfrak{u}}-\frac{1}{2}\log^{\mathfrak{u}}\operatorname{Li}_{1}^{\mathfrak{u}}$ is precisely the generator of $\mathcal{I}^{\operatorname{PL}}_{\{\ell\},2}$ studied in [DCW2], where $\ell$ denotes any prime. We note however that it is not an element of $\mathcal{I}^{\operatorname{PL}}_{S,2}$ for $|S|=2$.
@@ 1.3. Prior works
The present work is the first to tackle explicit Chabauty–Kim in depth >4>4 for $|S|\geq 2$. In this section, we give an overview of prior works on the subject.
The map loc was first made explicit in [DCW1] in depth $2$. In [BDCKW], the case $S=\emptyset$ in depths $1$ and $2$ was studied. In [DCW1], the equality $X(\mathbb{Z}_{S})=(X(\mathbb{Z}_{p})^{\mathrm{PL}}_{S,d})^{S_{3}}$ was verified for $S=\{2\}$ and $d=2$ with $p=3,5,7$, but turned out to fail with $p=11$. In [DCW2], Dan-Cohen and Wewers studied the case $S=\{2\}$ in depth $4$ and verified the equality $X(\mathbb{Z}_{S})=(X(\mathbb{Z}_{p})^{\mathrm{PL}}_{S,d})^{S_{3}}$ with $p=3,5,\ldots,29$ using two explicit $p$-adic Chabauty–Kim functions in depths $2$ and $4$ respectively [DCW2, Thm. 1.17.1]. The motivic framework of the method was further developed in [DC], resulting in an algorithm for determining $X(\mathbb{Z}_{S})$ whose halting is conditional on a list of conjectures. The polylogarithmic motivic Chabauty–Kim method for $X$ was fully developed in [CDC1], including the separation into the geometric and the arithmetic step. The geometric step was solved for $|S|=1$ in depth $4$, and the equality $X(\mathbb{Z}_{S})=(X(\mathbb{Z}_{p})^{\mathrm{PL}}_{S,d})^{S_{3}}$ was verified for $S=\{3\}$ in depth $4$ with $p=5,7$. It is also in [CDC1] that the necessity of symmetrising the polylogarithmic Chabauty–Kim locus via the $S_{3}$-action was discovered. In [CDC2], the authors describe an algorithm to determine the set $X(\mathbb{Z}_{S})$, based on the method of [CDC1], whose halting depends on a list of conjectures. Part of the work of the present paper can be seen as making their algorithm for the geometric step effective, a step that was perceived in [CDC1, CDC2] as simple but is far from it as soon as the size of $S$ or $d$ gets large. In related work, motivic Chabauty–Kim was recently developed for the surface $M_{0,5}$ in [DCJ] and was worked out explicitly for $S=\{2,3\}$ in depth $4$. A certain weight advantage encountered for $M_{0,5}$ allowed Dan-Cohen and the first author to work in depth $4$, a depth that is too small to handle $X$ when $|S|=2$. Their calculations solve in particular the arithmetic step for $X$ in depth $4$ for $S=\{2,3\}$. Recently, the case $S=\{2,3\}$ for $X$ was tackled in [martin] using a different method, namely the refined Chabauty–Kim method first developed in [bettsdogra]. This refined version of the method allows Lüdtke to work in depth $4$, resulting in new evidence for the refined Kim conjecture in this setting. The previous work [BB+] concerned the refined Chabauty–Kim method for $|S|=2$ in depth $2$. The recent work [BKL24] proves Kim’s conjecture for $S=\emptyset$ and the refined Kim conjecture for $S=\{2\}$, for all odd primes $p$.
@@ 1.4. Outline
Section˜2 contains preliminaries on mixed Tate motives and motivic fundamental groups. Section˜3 gives an overview of the polylogarithmic motivic Chabauty–Kim method. Section˜4 gives a description of the evaluation map (1.2) in terms of non-canonical coordinates of the Selmer scheme. This marks the end of the background material. Section˜5 contains the upper bound algorithm and the proofs of Theorem˜1.1(i)(i), $(iii)$. Finally, in Section˜6, we describe the resultant method, which culminates in the proof of Theorem˜1.1. Section˜7 contains a brief discussion of possible generalisations of the method to the case $|S|>2$.